POLYNOMIALS


FACTOR=

x3+x2-4x-4


1
Expert's answer
2021-03-16T07:11:15-0400
"x^3+x^2-4x-4=x^2(x+1)-4(x+1)"

"=(x+1)(x^2-4)=(x+1)(x+2)(x-2)"